# Linear combinations of latents in generative models: subspaces and beyond

### v7

@@ 1 Introduction
Generative models are a cornerstone of machine learning, with diverse applications including image synthesis, data augmentation, and creative content generation. Diffusion models (Ho et al., 2020; Song et al., 2020a; b) have emerged as a particularly effective approach to generative modeling for various modalities, such as for images (Ho et al., 2020), audio (Kong et al., 2020), video (Ho et al., 2022), and 3D models (Luo & Hu, 2021). A yet more recent approach to generative modelling is Flow Matching (Lipman et al., 2022), built upon Continuous Normalising Flows (Chen et al., 2018), that generalises diffusion to allow for different probability paths between data and latent distribution, e.g. defined through optimal transport (Gulrajani et al., 2017; Villani et al., 2009).
As well as generation, these models allow inversion where, by running the generative procedure in the opposite direction, data objects can be transformed deterministically into a corresponding realisation in the latent space. Such invertible connections between latent and data space provide a convenient mechanism for controlling generated objects by manipulating their latent vectors. The most common manipulation is to attempt semantically meaningful interpolation of two generated objects (Song et al., 2020a; b; Luo & Hu, 2021) by interpolating their corresponding latent vectors. However, the optimal choice of interpolant remains an open question, with simple approaches like linear interpolation leading to intermediates for which the model fails to generate plausible objects.
White (2016) argues that poor quality generation under linear interpolation is due to a mismatch between the norms of the intermediate vectors and those of the Gaussian vectors that the model has been trained to expect. Indeed, it is well-known that the squared norm of a $D$-dimensional unit Gaussian follows the chi-squared distribution. Consequently, likely samples are concentrated in a tight annulus around a radius $\sqrt{D}$ — a set which is not closed under linear interpolation. Because of this it is common to rely instead on spherical interpolation (Shoemake, 1985) (SLERP) that maintains similar norms as the endpoints. Motivated by poor performance of interpolation between the latent vectors provided by inverting natural images, alternatives to SLERP have recently been proposed (Samuel et al., 2023; Zheng et al., 2024). However, these procedures are costly, have parameters to tune, and — like SLERP — are challenging to generalise to other types of manipulations beyond interpolation, such as constructing the subspaces needed for latent space optimisation methods (Gómez-Bombarelli et al., 2018), or adapting to more diverse latent distributions beyond Gaussian.
@@
Figure 1: Low-dimensional latent subspaces. A 5-dimensional subspace from the flow matching model Stable Diffusion 3 (Esser et al., 2024) extracted using LOL (left) from the latents corresponding to images $\bm{x}_{1},\dots,\bm{x}_{5}$. The left plot show generations from uniform grid points across an axis-aligned slice of the subspace coordinate system, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of latents, which define basis vectors. The right plot shows the corresponding subspace without the proposed LOL transformation. See Figure 6, Figure 7 and Section G in the appendix for additional examples.
@@
Figure 2: Centroid determination. Generation using Stable Diffusion 2.1 (Rombach et al., 2022) from the centroid of the latents corresponding to images $\bm{x}_{1}$, $\bm{x}_{2}$, $\bm{x}_{3}$ using different methods. Note that our proposed method removes several artifacts, such as unrealistic headlights and chassi texture.
@@ 2.1 Generative Modelling with Latent Variables
Diffusion models learn a process that reverses the effect of gradually adding noise to training data (Ho et al., 2020). The noise level is governed by a schedule designed so that the noised samples at the final time index $T$ follow the latent distribution $\bm{x}^{(T)}\sim p(\bm{x})$. To sample from the diffusion model, one starts by drawing a sample from the latent distribution before iteratively evaluating the reverse process with learnt parameters $\bm{\theta}$, denoising step-by-step, generating a sequence $\{\bm{x}^{(T)},\bm{x}^{(T-1)},\ldots,\bm{x}^{(0)}\}$
@@
finally producing a generated object $\bm{x}^{(0)}$. Diffusion has also been extended to the continuous time setting by  Song et al. (2020b), where the diffusion is expressed as a stochastic differential equation. Note that, by using the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020a) or the probability flow formulation in Song et al. (2020b), the generation process can be made deterministic, i.e. the latent representation $\bm{x}^{(T)}\sim p(\bm{x})$ completely specifies the generated object $\bm{x}^{(0)}$.
Flow Matching Lipman et al. (2022) is an efficient approach to train Continuous Normalising Flows (CNF) Chen et al. (2018) — an alternative class of generative models that builds complicated distributions from simple latent distributions using differential equations. CNFs model the evolution of data points over continuous time using an ordinary differential equation (ODE) parameterized by a neural network, providing a deterministic relationship between latent and object space. Mathematically, the transformation of a latent sample $\bm{x}^{(T)}\sim p(\bm{x})$ at time $t=T$ to $\bm{x}^{(t)}$ at time $t$ is governed by $\bm{f}(\bm{x}^{(t)},t;\bm{\theta})$, where $\bm{f}(\cdot;\bm{\theta})$ is a neural network with parameters $\bm{\theta}$. Flow matching allows the transformation dynamics of the CNF to be learnt without needing to simulate the entire forward process during training, which improves stability and scalability.
@@ 2.2 Interpolation of Latents
Spherical Interpolation. However, as discussed in (White, 2016) in the context of Variational Autoencoders (Kingma, 2013) and Generative Adversarial Networks (Goodfellow et al., 2014), linear interpolation yields intermediates with unlikely norms for Gaussian samples and results in highly implausible generated objects (e.g. all-green images). Consequently, it is common to instead use spherical interpolation (SLERP) (Shoemake, 1985), which instead builds interpolants via
$\bm{y}_{\textrm{SLERP}}^{w}=\frac{\sin w\theta}{\sin\theta}\bm{x}_{1}+\frac{\sin((1-w)\theta)}{\sin\theta}\bm{x}_{2}\quad\textrm{for}\quad w\in[0,1]\quad\cos\theta=\frac{\langle\bm{x}_{1},\bm{x}_{2}\rangle}{||\bm{x}_{1}||||\bm{x}_{2}||},$
to maintain similar norms for the intermediates as the endpoints. SLERP has become popular and is the standard method for interpolation also in the more recent models (Song et al., 2020a; b).
Norm-aware Optimisation. Motivated by poor performance of SLERP when interpolating between the latent vectors corresponding to inverted natural images (rather than using those sampled directly from the model’s latent distribution),  Zheng et al. (2024) propose to add additional Gaussian noise to interpolants via the inversion procedure to control a trade-off between generation quality and adherence to the seed images (i.e. the images corresponding to the latents at the interpolation end-points). Alternatively, Samuel et al. (2023) advocate for Norm-Aware Optimisation (NAO) which uses back-propagation to identify interpolation paths $\gamma:[0,1]\rightarrow\mathds{R}^{d}$ that solve
@@
Figure 3: Likelihood and norm insufficient. Column (1) of each panel shows an image generated using a random sample from the associated Gaussian latent distribution for the diffusion model of Rombach et al. (2022) (left side) and the flow matching model of Esser et al. (2024) (right side). Columns (2) and (3) both show images generated from latents with the most likely norm according to their respective latent distribution. Columns (2) use the same Gaussian samples as in columns (1) but rescaled to have this norm, also yielding realistic images. Meanwhile, columns (3) show the failed generation from constant vectors $s\bm{I}$ scaled to have the most likely norm according to the latent distribution but lacking other characteristics (e.g. not having all-equal values) that the network was trained to expect, even though its likelihood $\mathcal{N}(s\bm{I};\bm{\mu},\bm{\Sigma})$ is typical of real samples. Moreover, the distribution mode $\bm{\mu}$, which also lacks needed characteristics , has vastly higher log likelihood than any realistic sample; -33875 and -135503 for the two models, respectively. See Table 2 for an example of failed generation using the mode $\bm{\mu}$.
@@ 3 Assessing validity of latents via distribution testing
The characteristics of the random variable $\bm{x}$ may be described by a collection of statistics, e.g. its mean, variance or norm. Above, we hypothesise that if a specified latent $\bm{x}_{\ast}$ has an extremely unlikely value for a characteristic for which the network has come to rely, then implausible generation will follow. A norm is unlikely if it has a low likelihood under its sampling distribution; for unit Gaussians it is $\chi(D)$, see Section 1. While the norm has been shown to often be an important statistic (Samuel et al., 2023), our Figure 3 illustrates on a popular diffusion model (Rombach et al., 2022) and a flow matching model (Esser et al., 2024) that a latent with a likely norm — even the most likely — can still induce failure in the generative process if other evidently critical characteristics are unmet.
@@
Figure 4: Normality testing of latent vectors obtained from inversion. We show the LPIPS (Zhang et al., 2018) reconstruction errors (second row, in red) of 200 inverted randomly selected images across 50 random classes from ImageNet1k (Deng et al., 2009), the p-values of their inversions (third row), and rejection rates (bottom row) of the Kolmogorov-Smirnov normality test applied to the corresponding latent obtained from inversion under various step budgets. We use the diffusion model of Rombach et al. (2022), always using its maximum number of steps (999) for generation, and denote the 10th, 50th and 90th percentiles with black lines. The first row shows image reconstructions using its inversion at each budget, highlighted in red when the latent was rejected ($p<1e^{-3}$), with the interpretation that the characteristics of the latent were unlikely for a real sample according to the KS test. We note the strong correlation between inversion budgets providing low reconstruction errors and those for which the p-values of the latents are realistic — taking values likely to occur by chance for real samples. However, as we will see in Figure 5, there are still many latents with low reconstruction error yet extremely low p-value, and this often severely affects the quality of its interpolants.
In general, neural networks have many degrees of freedom, and so it is difficult to determine and list the input characteristics that a specific model relies on. Therefore, rather than trying to determine all necessary statistics for each model, we propose instead to rely on standard methods for distribution testing, a classic area of statistics  (Razali et al., 2011; Kolmogorov, 1933; Shapiro & Wilk, 1965) — here exploring the (null) hypothesis that a latent vector $\bm{x}_{\ast}\in\mathbb{R}^{D}$ is drawn from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. Popular normality tests consider broad statistics associated with normality. For example, the Kolmogorov–Smirnov (Kolmogorov, 1933) test considers the distribution of the largest absolute difference between the empirical and the theoretical cumulative density function across all values.
@@ 3.1 The sample characteristics of seed latents affect their interpolations
Figure 5: Lack of normality is linked to failure of interpolants The left and middle panels shows LPIPS (Zhang et al., 2018) reconstruction error (after 999 generation steps) and the Kolmogorov-Smirnov p-value for all inversions presented in Figure 4, split into two plots due to the vast dynamic range of p-values. Although latents with high (realistic) p-values tend to have low reconstruction errors, there are many latents with low reconstruction errors that also have low p-values. The right panel shows Q-Align visual quality scores (Wu et al., 2023) for spherical interpolants between pairs of inversions selected from matching ImageNet1k (Deng et al., 2009) image classes, demonstrating that choosing seed latents with both low reconstruction error ($<0.05$) and high p-values ($>1e^{-3}$) allows us to avoid low-quality interpolants that would arise when choosing seeds by reconstruction error alone. For reference, we include examples of interpolants at each visual quality level.
@@ 4 Linear combinations of latents
where $w_{k}\in\mathbb{R}$, $\bm{w}=[w_{1},w_{2},\dots,w_{K}]$ and $\bm{X}=[\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{K}]$. Then we have that $\bm{y}$ is also a Gaussian random variable, with mean and covariance
@@ 5.1 Interpolation and centroid determination
For the application of interpolation, we compare our proposed LOL to linear interpolation (LERP), spherical linear interpolation (SLERP), and Norm-Aware Optimization (NAO) Samuel et al. (2023).
In contrast to the all the other considered interpolation methods (including LOL) which only involve closed-form expressions, NAO requires numerical optimisation. We closely follow the evaluation protocol in Samuel et al. (2023), basing the experiments on Stable Diffusion (SD) 2.1 (Rombach et al., 2022) and inversions of random images from 50 random classes from ImageNet1k (Deng et al., 2009), and assess visual quality and preservation of semantics using Fréchet Inception Distance (FID) (Heusel et al., 2017) and class prediction accuracy, respectively. For the interpolation we (randomly without replacement) pair the 50 images per class into 25 pairs, forming 1250 image pairs in total.
@@
LERP $3.92$% $198.8888$ $6e^{-3}$s
SLERP $64.56$% $42.5827$ $9e^{-3}$s
NAO 62.133 333 333 333 3362.1333333333333362.133\,333\,333\,333\,3362.133 333 333 333 33% $46.0230$ $30$s
LOL (ours) 67.386 666 666 666 6667.3866666666666667.386\,666\,666\,666\,6667.386 666 666 666 66% $38.8714$ $6e^{-3}$s
@@
Euclidean 0.285 714 285 714 285 70.28571428571428570.285\,714\,285\,714\,285\,70.285 714 285 714 285 7% 310.415 342 373 754 75310.41534237375475310.415\,342\,373\,754\,75310.415 342 373 754 75 $4e^{-4}$s
Standardised Euclidean 44.571 428 571 428 57344.57142857142857344.571\,428\,571\,428\,57344.571 428 571 428 573% 88.839 964 830 394 788.839964830394788.839\,964\,830\,394\,788.839 964 830 394 7 $1e^{-3}$s
Mode norm Euclidean 44.571 428 571 428 57344.57142857142857344.571\,428\,571\,428\,57344.571 428 571 428 573% 88.377 760 874 495 3888.3777608744953888.377\,760\,874\,495\,3888.377 760 874 495 38 $1e^{-3}$s
NAO $44.000$% 92.967 278 310 416 1392.9672783104161392.967\,278\,310\,416\,1392.967 278 310 416 13 $90$s
LOL (ours) 46.285 714 285 714 28646.28571428571428646.285\,714\,285\,714\,28646.285 714 285 714 286% 87.713 436 033 613 3387.7134360336133387.713\,436\,033\,613\,3387.713 436 033 613 33 $6e^{-4}$s
@@
In Table 1 we show that our method outperforms or maintains the performance of the baselines in terms of FID distance and accuracy, as calculated using a pre-trained classifier following the evaluation methodology of Samuel et al. (2023). For an illustration of centroids and interpolations, see Figure 2 and Figure 16, respectively. The evaluation time of an interpolation path and centroid, shown with one digit of precision, illustrate that the closed-form expressions are significantly faster than NAO. Surprisingly, NAO did not perform as well as spherical interpolation and several other baselines, despite using their implementation, which was outperforming these methods in Samuel et al. (2023). We note one discrepancy is that we report FID distances using (2048) high-level features, while in their work they are using (64) low-level features, which in Seitzer (2020) is recommended against as it does not necessarily correlate with visual quality. In the appendix we include FID distances using all settings of features. We note that, in our setup, the baselines perform substantially better than reported in Samuel et al. (2023) — including NAO in terms of FID distances using the 64 low-level feature setting (1.30 in our setup vs 6.78 in their setup), and class accuracy during interpolation (62% vs 52%). See Section H in the appendix for more details and ablations.
@@
Figure 6: Low-dimensional subspaces. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from the flow matching model Stable Diffusion 3 (Esser et al., 2024) over uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,2\}$ and $\{3,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors, which through LOL all yield high-quality generations.

### v8

@@ 1 Introduction
Generative models are a cornerstone of machine learning, with diverse applications including image synthesis, data augmentation, and creative content generation. Diffusion models (ho2020denoising; song2020denoising; song2020score) have emerged as a particularly effective approach to generative modeling for various modalities, such as for images (ho2020denoising), audio (kong2020diffwave), video (ho2022video), and 3D models (luo2021diffusion). A yet more recent approach to generative modelling is Flow Matching (lipman2022flow), built upon Continuous Normalising Flows (chen2018neural), that generalises diffusion to allow for different probability paths between data and latent distribution, e.g. defined through optimal transport (gulrajani2017improved; villani2009optimal).
As well as generation, these models allow inversion where, by running the generative procedure in the opposite direction, data objects can be transformed deterministically into a corresponding realisation in the latent space. Such invertible connections between latent and data space provide a convenient mechanism for controlling generated objects by manipulating their latent vectors. The most common manipulation is to attempt semantically meaningful interpolation of two generated objects (song2020denoising; song2020score; luo2021diffusion) by interpolating their corresponding latent vectors. However, the optimal choice of interpolant remains an open question, with simple approaches like linear interpolation leading to intermediates for which the model fails to generate plausible objects.
white2016sampling argues that poor quality generation under linear interpolation is due to a mismatch between the norms of the intermediate vectors and those of the Gaussian vectors that the model has been trained to expect. Indeed, it is well-known that the squared norm of a $D$-dimensional unit Gaussian follows the chi-squared distribution. Consequently, likely samples are concentrated in a tight annulus around a radius $\sqrt{D}$ — a set which is not closed under linear interpolation. Because of this it is common to rely instead on spherical interpolation (shoemake1985animating) (SLERP) that maintains similar norms as the endpoints. Motivated by poor performance of interpolation between the latent vectors provided by inverting natural images, alternatives to SLERP have recently been proposed (samuel2024norm; zheng2024noisediffusion). However, these procedures are costly, have parameters to tune, and — like SLERP — are challenging to generalise to other types of manipulations beyond interpolation, such as constructing the subspaces needed for latent space optimisation methods (gomez2018automatic), or adapting to more diverse latent distributions beyond Gaussian.
@@
Figure 1: Low-dimensional latent subspaces. A 5-dimensional subspace from the flow matching model Stable Diffusion 3 (esser2024scaling) extracted using LOL (left) from the latents corresponding to images $\bm{x}_{1},\dots,\bm{x}_{5}$. The left plot show generations from uniform grid points across an axis-aligned slice of the subspace coordinate system, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of latents, which define basis vectors. The right plot shows the corresponding subspace without the proposed LOL transformation. See Figure 6, Figure 7 and Section G in the appendix for additional examples.
@@
Figure 2: Centroid determination. Generation using Stable Diffusion 2.1 (rombach2022high) from the centroid of the latents corresponding to images $\bm{x}_{1}$, $\bm{x}_{2}$, $\bm{x}_{3}$ using different methods. Note that our proposed method removes several artifacts, such as unrealistic headlights and chassi texture.
@@ 2.1 Generative Modelling with Latent Variables
Diffusion models learn a process that reverses the effect of gradually adding noise to training data (ho2020denoising). The noise level is governed by a schedule designed so that the noised samples at the final time index $T$ follow the latent distribution $\bm{x}^{(T)}\sim p(\bm{x})$. To sample from the diffusion model, one starts by drawing a sample from the latent distribution before iteratively evaluating the reverse process with learnt parameters $\bm{\theta}$, denoising step-by-step, generating a sequence $\{\bm{x}^{(T)},\bm{x}^{(T-1)},\ldots,\bm{x}^{(0)}\}$
@@
finally producing a generated object $\bm{x}^{(0)}$. Diffusion has also been extended to the continuous time setting by  song2020score, where the diffusion is expressed as a stochastic differential equation. Note that, by using the Denoising Diffusion Implicit Model (DDIM) (song2020denoising) or the probability flow formulation in song2020score, the generation process can be made deterministic, i.e. the latent representation $\bm{x}^{(T)}\sim p(\bm{x})$ completely specifies the generated object $\bm{x}^{(0)}$.
Flow Matching lipman2022flow is an efficient approach to train Continuous Normalising Flows (CNF) chen2018neural — an alternative class of generative models that builds complicated distributions from simple latent distributions using differential equations. CNFs model the evolution of data points over continuous time using an ordinary differential equation (ODE) parameterized by a neural network, providing a deterministic relationship between latent and object space. Mathematically, the transformation of a latent sample $\bm{x}^{(T)}\sim p(\bm{x})$ at time $t=T$ to $\bm{x}^{(t)}$ at time $t$ is governed by $\bm{f}(\bm{x}^{(t)},t;\bm{\theta})$, where $\bm{f}(\cdot;\bm{\theta})$ is a neural network with parameters $\bm{\theta}$. Flow matching allows the transformation dynamics of the CNF to be learnt without needing to simulate the entire forward process during training, which improves stability and scalability.
@@ 2.2 Interpolation of Latents
Spherical Interpolation. However, as discussed in (white2016sampling) in the context of Variational Autoencoders (kingma2013auto) and Generative Adversarial Networks (goodfellow2014generative), linear interpolation yields intermediates with unlikely norms for Gaussian samples and results in highly implausible generated objects (e.g. all-green images). Consequently, it is common to instead use spherical interpolation (SLERP) (shoemake1985animating), which instead builds interpolants via
@@
to maintain similar norms for the intermediates as the endpoints. SLERP has become popular and is the standard method for interpolation also in the more recent models (song2020denoising; song2020score).
Norm-aware Optimisation. Motivated by poor performance of SLERP when interpolating between the latent vectors corresponding to inverted natural images (rather than using those sampled directly from the model’s latent distribution),  zheng2024noisediffusion propose to add additional Gaussian noise to interpolants via the inversion procedure to control a trade-off between generation quality and adherence to the seed images (i.e. the images corresponding to the latents at the interpolation end-points). Alternatively, samuel2024norm advocate for Norm-Aware Optimisation (NAO) which uses back-propagation to identify interpolation paths $\gamma:[0,1]\rightarrow\mathds{R}^{d}$ that solve
$\inf_{\gamma}-\int\log\mathds{P}(\gamma(s))ds\quad\textrm{s.t.}\quad\gamma(0)=\bm{x}_{1},\gamma(1)=\bm{x}_{2},$
@@
Figure 3: Likelihood and norm insufficient. Column (1) of each panel shows an image generated using a random sample from the associated Gaussian latent distribution for the diffusion model of rombach2022high (left side) and the flow matching model of esser2024scaling (right side). Columns (2) and (3) both show images generated from latents with the most likely norm according to their respective latent distribution. Columns (2) use the same Gaussian samples as in columns (1) but rescaled to have this norm, also yielding realistic images. Meanwhile, columns (3) show the failed generation from constant vectors $s\bm{I}$ scaled to have the most likely norm according to the latent distribution but lacking other characteristics (e.g. not having all-equal values) that the network was trained to expect, even though its likelihood $\mathcal{N}(s\bm{I};\bm{\mu},\bm{\Sigma})$ is typical of real samples. Moreover, the distribution mode $\bm{\mu}$, which also lacks needed characteristics , has vastly higher log likelihood than any realistic sample; -33875 and -135503 for the two models, respectively. See Table 2 for an example of failed generation using the mode $\bm{\mu}$.
@@ 3 Assessing validity of latents via distribution testing
The characteristics of the random variable $\bm{x}$ may be described by a collection of statistics, e.g. its mean, variance or norm. Above, we hypothesise that if a specified latent $\bm{x}_{\ast}$ has an extremely unlikely value for a characteristic for which the network has come to rely, then implausible generation will follow. A norm is unlikely if it has a low likelihood under its sampling distribution; for unit Gaussians it is $\chi(D)$, see Section 1. While the norm has been shown to often be an important statistic (samuel2024norm), our Figure 3 illustrates on a popular diffusion model (rombach2022high) and a flow matching model (esser2024scaling) that a latent with a likely norm — even the most likely — can still induce failure in the generative process if other evidently critical characteristics are unmet.
@@
Figure 4: Normality testing of latent vectors obtained from inversion. We show the LPIPS (zhang2018unreasonable) reconstruction errors (second row, in red) of 200 inverted randomly selected images across 50 random classes from ImageNet1k (deng2009imagenet), the p-values of their inversions (third row), and rejection rates (bottom row) of the Kolmogorov-Smirnov normality test applied to the corresponding latent obtained from inversion under various step budgets. We use the diffusion model of rombach2022high, always using its maximum number of steps (999) for generation, and denote the 10th, 50th and 90th percentiles with black lines. The first row shows image reconstructions using its inversion at each budget, highlighted in red when the latent was rejected ($p<1e^{-3}$), with the interpretation that the characteristics of the latent were unlikely for a real sample according to the KS test. We note the strong correlation between inversion budgets providing low reconstruction errors and those for which the p-values of the latents are realistic — taking values likely to occur by chance for real samples. However, as we will see in Figure 5, there are still many latents with low reconstruction error yet extremely low p-value, and this often severely affects the quality of its interpolants.
In general, neural networks have many degrees of freedom, and so it is difficult to determine and list the input characteristics that a specific model relies on. Therefore, rather than trying to determine all necessary statistics for each model, we propose instead to rely on standard methods for distribution testing, a classic area of statistics  (razali2011power; an1933sulla; shapiro1965analysis) — here exploring the (null) hypothesis that a latent vector $\bm{x}_{\ast}\in\mathbb{R}^{D}$ is drawn from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. Popular normality tests consider broad statistics associated with normality. For example, the Kolmogorov–Smirnov (an1933sulla) test considers the distribution of the largest absolute difference between the empirical and the theoretical cumulative density function across all values.
@@ 3.1 The sample characteristics of seed latents affect their interpolations
Figure 5: Lack of normality is linked to failure of interpolants The left and middle panels shows LPIPS (zhang2018unreasonable) reconstruction error (after 999 generation steps) and the Kolmogorov-Smirnov p-value for all inversions presented in Figure 4, split into two plots due to the vast dynamic range of p-values. Although latents with high (realistic) p-values tend to have low reconstruction errors, there are many latents with low reconstruction errors that also have low p-values. The right panel shows Q-Align visual quality scores (wu2023q) for spherical interpolants between pairs of inversions selected from matching ImageNet1k (deng2009imagenet) image classes, demonstrating that choosing seed latents with both low reconstruction error (<0.05<0.05) and high p-values ($1e^{-3}$) allows us to avoid low-quality interpolants that would arise when choosing seeds by reconstruction error alone. For reference, we include examples of interpolants at each visual quality level.
@@ 4 Linear combinations of latents
[…]
@@ 5.1 Interpolation and centroid determination
For the application of interpolation, we compare our proposed LOL to linear interpolation (LERP), spherical linear interpolation (SLERP), and Norm-Aware Optimization (NAO) samuel2024norm.
In contrast to the all the other considered interpolation methods (including LOL) which only involve closed-form expressions, NAO requires numerical optimisation. We closely follow the evaluation protocol in samuel2024norm, basing the experiments on Stable Diffusion (SD) 2.1 (rombach2022high) and inversions of random images from 50 random classes from ImageNet1k (deng2009imagenet), and assess visual quality and preservation of semantics using Fréchet Inception Distance (FID) (heusel2017gans) and class prediction accuracy, respectively. For the interpolation we (randomly without replacement) pair the 50 images per class into 25 pairs, forming 1250 image pairs in total.
@@
LERP 3.923.92% $198.8888$ $6e^{-3}$s
SLERP 64.5664.56% $42.5827$ $9e^{-3}$s
NAO 62.133 333 333 333 3362.133\,333\,333\,333\,33% $46.0230$ $30$s
LOL (ours) 67.386 666 666 666 6667.386\,666\,666\,666\,66% $38.8714$ $6e^{-3}$s
@@
Euclidean 0.285 714 285 714 285 70.285\,714\,285\,714\,285\,7% 310.415 342 373 754 75310.415\,342\,373\,754\,75 $4e^{-4}$s
Standardised Euclidean 44.571 428 571 428 57344.571\,428\,571\,428\,573% 88.839 964 830 394 788.839\,964\,830\,394\,7 $1e^{-3}$s
Mode norm Euclidean 44.571 428 571 428 57344.571\,428\,571\,428\,573% 88.377 760 874 495 3888.377\,760\,874\,495\,38 $1e^{-3}$s
NAO 44.00044.000% 92.967 278 310 416 1392.967\,278\,310\,416\,13 $90$s
LOL (ours) 46.285 714 285 714 28646.285\,714\,285\,714\,286% 87.713 436 033 613 3387.713\,436\,033\,613\,33 $6e^{-4}$s
@@
In Table 1 we show that our method outperforms or maintains the performance of the baselines in terms of FID distance and accuracy, as calculated using a pre-trained classifier following the evaluation methodology of samuel2024norm. For an illustration of centroids and interpolations, see Figure 2 and Figure 16, respectively. The evaluation time of an interpolation path and centroid, shown with one digit of precision, illustrate that the closed-form expressions are significantly faster than NAO. Surprisingly, NAO did not perform as well as spherical interpolation and several other baselines, despite using their implementation, which was outperforming these methods in samuel2024norm. We note one discrepancy is that we report FID distances using (2048) high-level features, while in their work they are using (64) low-level features, which in Seitzer2020FID is recommended against as it does not necessarily correlate with visual quality. In the appendix we include FID distances using all settings of features. We note that, in our setup, the baselines perform substantially better than reported in samuel2024norm — including NAO in terms of FID distances using the 64 low-level feature setting (1.30 in our setup vs 6.78 in their setup), and class accuracy during interpolation (62% vs 52%). See Section H in the appendix for more details and ablations.
@@
Figure 6: Low-dimensional subspaces. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from the flow matching model Stable Diffusion 3 (esser2024scaling) over uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,2\}$ and $\{3,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors, which through LOL all yield high-quality generations.